# Benchmarking Vision, Language, & Action Models on Robotic Learning Tasks

### v1

@@ 1 Introduction
The development of generalizable robotic systems remains a central challenge in machine learning and robotics. While recent advances have enabled impressive capabilities in controlled settings, learned policies frequently fail to generalize beyond their training distribution. This limitation manifests across multiple dimensions: models struggle to respond appropriately to novel task instructions [4, 23], handle variations in object positions and orientations [3], adapt to changing lighting conditions or partially obscured scenes [6], interact with previously unseen objects, or maintain performance in the presence of distracting objects [25, 22]. This brittleness presents a significant barrier to deploying learned robotics systems in unconstrained real-world environments.
Concurrent advances in foundation models, particularly in vision and language domains, offer a compelling potential solution to these challenges. Web-scale training has enabled models to realize SotA capabilities in visual recognition [15, 17], complex reasoning about object-agent interactions [2, 8, 24], code generation [5], and multimodal understanding. These models exhibit precisely the kinds of robust generalization capabilities that traditional robotics approaches have found challenging to achieve. The semantic reasoning, problem-solving, and visual interpretation capabilities demonstrated by these models would be tremendously valuable for developing generalist robots capable of performing diverse tasks in dynamic real-world environments.
This direction aligns with an emerging trend in machine learning regarding the advantages of unified neural sequence models. Such models continue to show performance improvements even at the frontier of data, compute, and model scale [11, 10]. Moreover, historical trends suggest that generic models capable of effectively leveraging computation tend to eventually supersede specialized domain-specific approaches [21]. Unified sequence models offer several compelling advantages: they eliminate the need for hand-crafted policy architectures with domain-specific inductive biases, can leverage diverse training data through sequence serialization, and demonstrate consistent improvements with increased scale.
However, significant challenges remain in applying these models to robotics. While foundation models are typically trained on billions of tokens and images from the web, collecting comparable quantities of robotic interaction data remains infeasible in the near term [7, 13]. Additionally, the direct application of foundation models to robotics tasks presents fundamental technical challenges: these models excel at semantic reasoning and high-level understanding, but robots require precise, grounded, low-level control actions, such as Cartesian end-effector commands. Recent work has attempted to bridge this gap by incorporating language models (LLMs) and vision-language models (VLMs) into robotics systems (Ahn et al., 2022; Driess et al., 2023; Vemprala et al., 2023). However, many of these approaches restrict foundation models to high-level planning, effectively using them as sophisticated state machines that parse commands into primitive actions, which are then executed by separate low-level controllers that cannot leverage the rich semantic knowledge encoded in the foundation models.
More recent research has explored bootstrapping robotics representations using pretrained language and vision-language models [19, 18, 12]. Applying these modules in components in planning systems has also been explored [8, 20]. A promising direction has emerged in the form of vision-language-action models (VLAs), which often involve extending different kinds of pretrained foundation models for robotics by pretraining [3], cotraining [22] or fine-tuning [14, 4, 16] visually-conditioned language models to control actions for robots. These models have demonstrated notable transfer to new task scenarios, a compelling first step on the path towards generally useful policies for robots and agents.
@@
  • Detailed profiling results for an initial set of VLM, VLA, and emerging "generalist" models, providing insights into their capabilities and limitations
  • Analysis of generalization
  • A systematic set of evaluation splits and metrics specifically designed for robotics learning tasks in the widely-used OpenX Dataset
  • A general framework for mapping VLMs to other modality classes, with particular emphasis on action spaces
  • Open-source software infrastructure for downloading, managing, and utilizing the benchmark data
Through this work, we aim to provide the robotics learning community with robust tools and methodologies for assessing and comparing these emerging approaches, facilitating progress in this rapidly evolving field and helping to bridge the gap between foundation models and practical robotics applications. Importantly, this is the first foray into a new large scale generalist action model benchmark, which we discuss in the context of Future Work.
@@ 2 Related Work
MultiBench represents one of the first systematic attempts to evaluate multimodal learning across diverse domains, spanning healthcare, robotics, affective computing, and finance. Similar to our work, MultiBench emphasizes the importance of evaluating multiple aspects of model performance, including generalization, complexity, and robustness. However, while MultiBench covers a broad range of domains, its robotics evaluation is limited in scope. MMMU provides another comprehensive benchmark focused on college-level multimodal understanding. The authors evaluate models across technical disciplines like engineering and science through expert-level problems requiring nuanced perception and domain-specific knowledge, but do not specifically address robotics control tasks.
@@
The evolution of multimodal evaluation has progressed from single-task benchmarks like VQA, OK-VQA, MSCOCO, and GQA to more comprehensive evaluation frameworks. Recent benchmarks span various capabilities, from basic OCR to adversarial robustness and hallucination detection (e.g., POPE and HaELM). More holistic evaluations have emerged through benchmarks like LAMM, LVLM-eHub, SEED, MMBench, and MM-Vet. Specialized benchmarks such as MathVista focus on specific domains like mathematical reasoning, while GAIA tests fundamental abilities in reasoning and multimodality handling.
@@
The evolution of robotics datasets has demonstrated considerable diversity across various dimensions, particularly with the advancement of imitation learning and behavior cloning (BC). While many robotics benchmarks focus on evaluating model adaptability to new tasks, functionalities, or environments, there remains a gap in systematically evaluating different BC models at scale in both simulated and real-world settings. THE COLOSSEUM addresses this gap by providing a systematic evaluation framework focused on robotic manipulation, evaluating generalization across 14 different environmental perturbations. Similar efforts include FactorWorld, which examines 11 variation factors across 19 tasks, and KitchenShift, which evaluates zero-shot generalization across 7 variation factors in kitchen environments.Several other specialized robotics benchmarks have emerged: RLBench offers a suite of 100 manipulation tasks in simulation; RAVENS focuses on vision-based manipulation; and FurnitureBench provides reproducible real-world benchmarks for long-horizon complex manipulation. LIBERO offers benchmarks for knowledge transfer in lifelong robot learning, while FMB (Functional Manipulation Benchmark) emphasizes generalizable robotic learning across complex tasks. Recent work has also introduced DUDE for robotic document manipulation and ProcTHOR for large-scale embodied AI using procedural generation.
@@ 3.1 Data
For version 0.1 of our benchmark, we utilize 53 of the 72 available OpenX datasets, as detailed in Figure [X]. We present results for 20 of these datasets for all 3 models, and have the full 53 for JAT.This subset was selected to ensure comprehensive coverage across different task types, embodiments, and environmental conditions while maintaining data quality and consistency. For datasets that did not include pre-defined evaluation sets, we have created and provided new evaluation splits to ensure robust assessment of model performance. The training splits of these 53 datasets comprise approximately 32 terabytes of data.
@@
The complete list of included datasets and their characteristics is provided in the appendix, along with details about our evaluation split creation methodology.
@@ 3.1.1 Dataset Curation
Our curation methodology consisted of several steps. First, we conducted a high-level review of dataset quality and accessibility, which resulted in the exclusion of three datasets: Austin BUDS, Austin Sailor, and Stanford Kuka Multimodal. For datasets that contained only training splits, we performed a detailed comparative analysis based on the robot platform used for data collection. This analysis considered multiple features: Robot model and morphology, Gripper specifications, Action space characteristics, Sensor configuration (number and type of RGB cameras, depth cameras, and wrist-mounted cameras), Presence of language annotations, Availability of camera calibration data, Inclusion of proprioceptive information
@@
Several additional datasets were excluded from version 0.1 of our benchmark due to technical limitations in their accessibility through the TensorFlow Datasets (TFDS) builder, which is the recommended data loading mechanism for OpenX. These compatibility issues will be addressed in future versions of the benchmark as the underlying infrastructure evolves. This careful curation process results in a benchmark that balances comprehensive coverage with practical considerations, ensuring that the included datasets provide meaningful evaluation scenarios while maintaining manageable computational requirements. The complete list of included and excluded datasets, along with the specific rationale for each curation decision, is provided in Appendix [X].
@@ 3.2 Models
In our evaluation, we focus on three recent vision-language-action (VLA) models that represent the current state-of-the-art in generalist robot learning: JAT (Jack of All Trades), GPT-4o, and OpenVLA. These models are particularly noteworthy for their ability to handle multiple modalities and their demonstrated capabilities in robotic control tasks.
@@
JAT emphasizes broad "generalist" multi-modal capabilities GPT-4o focuses on omni-modal processing with robust safety considerations OpenVLA prioritizes open-source accessibility while maintaining competitive performance with larger closed-source models
@@ 3.3 Evaluation Metrics
We report several variations of MSE to provide comprehensive performance analysis:
@@
    Average MSE (AMSE): Computed as the mean MSE across all trajectories in a dataset, AMSE enables direct comparison of model performance across different datasets and architectures.
@@
    Normalized AMSE (NAMSE): Calculated as $(prediction-min\_prediction)/(max\_prediction-min\_prediction)$, this metric normalizes predictions to each model’s prediction range, facilitating more equitable cross-dataset for a single model comparisons by accounting for different scales in model outputs.
@@
    Completion Rate: We assess successful completion by comparing final predicted actions with ground truth final actions. While this serves as an approximate measure of task completion, it provides valuable insights into models’ ability to reach target states across trajectories.
The combination of these metrics allows us to evaluate both the fine-grained accuracy of action predictions and the overall task completion capabilities of different models. This is particularly important in offline robotics, where environments and rewards are not available for policy evaluation.
@@ 4.1 Profiling Configuration
  • Multi-Image Processing: We provide all images corresponding to multiple views to GPT.
@@
    • Continuous normalization: For $[0,1]$ to $[-1,1]$ conversion, we apply the formula: $y=2\cdot(x-orig_{low})/(orig_{high}-orig_{low})-1$. This was used by the authors in [14].
@@
Common Considerations
Across all models, we encountered cases where image observations were unavailable due to non-standard image key naming (e.g., ‘agentview_rgb’, ‘frontright_fisheye_image’) in some datasets. These edge cases were handled appropriately for each model while maintaining consistent evaluation protocols.
@@ 5 Results & Discussion
Figure 1: AMSE Across All Datasets
Table 1: Dataset Coverage and Action Space Characteristics for OpenVLA Evaluation
Dataset Name Registered Dataset Name In Pretraining Action Space Type
@@ 5.1 Average Model Performance Analysis
  • PR2 robot tasks and kitchen manipulation scenarios consistently show high error rates
  • Complex manipulation tasks, particularly those involving large movements or multi-step sequences
  • Tasks requiring significant temporal reasoning or complex action sequences
@@ 5.1.2 Implications for Future Development
Figure 2: Normalized AMSE For GPT4o
Figure 3: Normalized AMSE For JAT
Figure 4: Normalized AMSE For OpenVLA
@@ 5.2.1 Model-Specific Performance Patterns
  • Exceptionally strong performance on certain tasks (e.g., toto with NAMSE: 0.003)
@@
  • Notable performance spike on ucsd_kitchen_dataset (NAMSE $\sim$0.57)
@@ 5.2.2 Cross-Model Insights
  • Kitchen manipulation tasks and complex multi-step operations consistently show higher NAMSE

### v2

@@ Abstract
Figure 1: Multinet v0.1 overview. Benchmarking a SoTA VLM, SoTA VLA, and novel generalist model on 20 real-world robotics datasets by comparing the action predicted by the models, with the ground truth action from the dataset, at every timestep.
@@ 1 Introduction
The quest for robust, generalizable robotic systems continues to pose a fundamental challenge in machine learning and robotics research. Despite significant progress in controlled environments, current systems exhibit limited generalization beyond their training conditions. These limitations span numerous dimensions: systems fail when encountering unfamiliar task descriptions [6, 42], struggle with spatial variations in object configurations [5], perform poorly under variable lighting or occlusion [8], and show degraded performance when interacting with novel objects or in cluttered environments [45, 39]. These generalization challenges significantly hinder the deployment of learned robotic systems in unconstrained environments.
Recent breakthroughs in foundation models, especially in vision and language processing, suggest a promising path forward. These models, trained on web-scale datasets, have achieved remarkable capabilities in visual understanding [23, 33], sophisticated reasoning about interactions between objects and agents [3, 11, 43], software development [7], and cross-modal comprehension. The robust generalization exhibited by these models addresses precisely the challenges that have historically limited robotics systems. Their advanced capabilities in semantic understanding, problem-solving, and visual processing could revolutionize the development of versatile robots capable of handling diverse tasks in dynamic environments.
This approach corresponds with a broader trend in machine learning toward unified neural sequence architectures. These models demonstrate continued performance gains at the boundaries of data volume, computational resources, and model complexity [19, 15]. This pattern aligns with historical observations suggesting that general-purpose models efficiently utilizing computational resources tend to outperform specialized solutions [38]. The advantages of unified sequence models are multifaceted: they remove the requirement for custom policy architectures with domain-specific assumptions, enable the use of diverse training data through sequence-based representation, and show reliable improvements with increasing scale.
Nevertheless, adapting these models for robotics applications presents substantial challenges. The vast scale of training data available for foundation models - billions of tokens and images from the internet - far exceeds what is currently feasible to collect for robot interactions [9, 21]. Moreover, while foundation models excel at abstract reasoning and high-level comprehension, robotic control requires precise, physically grounded actions, such as specific end-effector movements. Recent research has explored integrating language models (LLMs) and vision-language models (VLMs) into robotics frameworks [2, 11, 41]. However, many current approaches limit foundation models to high-level planning roles, using them essentially as advanced state machines that convert commands into basic actions, executed by separate low-level controllers unable to access the models’ rich semantic understanding.
Current research initiatives have investigated leveraging pretrained language and vision-language models to enhance robotic representations [36, 34, 20]. These components have also been integrated into planning systems [11, 37]. A particularly promising development has been the emergence of vision-language-action models (VLAs), which extend foundation models for robotics through various approaches including pretraining [5, 39] or fine-tuning [22, 6, 30]. These models have shown encouraging results in transferring to novel tasks, marking an important advancement toward developing generally capable robotic systems.
@@
  • Detailed profiling results for an initial set of VLM, VLA, and emerging “generalist” models, providing insights into their capabilities and limitations.
  • Analysis of model generalization to a diverse set of real-world robotics datasets comprising a wide variety of tasks and environments.
  • A systematic set of evaluation splits and metrics specifically designed for robotics learning tasks in the widely-used OpenX Dataset.
  • A general framework for mapping VLMs to other modality classes, with particular emphasis on action spaces.
  • Open-source software infrastructure for downloading, managing, and utilizing the benchmark data.
Through this work, we aim to provide the robotics learning community with robust tools and methodologies for assessing and comparing these emerging approaches, facilitating progress in this rapidly evolving field and helping to bridge the gap between foundation models and practical robotics applications. Importantly, this is the first foray into a new large-scale generalist action model benchmark, called MultiNet v0.1, which we discuss in the context of Future Work.
@@ 2 Related Work
MultiBench [25] represents one of the first systematic attempts to evaluate multimodal learning across diverse domains, spanning healthcare, robotics, affective computing, and finance. Similar to our work, MultiBench emphasizes the importance of evaluating multiple aspects of model performance, including generalization, complexity, and robustness. However, while MultiBench covers a broad range of domains, its robotics evaluation is limited in scope. MMMU [50] provides another comprehensive benchmark focused on college-level multimodal understanding. The authors evaluate models across technical disciplines like engineering and science through expert-level problems requiring nuanced perception and domain-specific knowledge, but do not specifically address robotics control tasks.
@@
The evolution of multimodal evaluation has progressed from single-task benchmarks like VQA [4], OK-VQA [31], MSCOCO [26], and GQA [17] to more comprehensive evaluation frameworks. Recent benchmarks span various capabilities, from basic OCR to adversarial robustness and hallucination detection (e.g., POPE [24] and HaELM [44]). More holistic evaluations have emerged through benchmarks like LAMM [48], LVLM-eHub [47], SEED [13], MMBench [51], and MM-Vet [49]. Specialized benchmarks such as MathVista [28] focus on specific domains like mathematical reasoning, while GAIA [32] tests fundamental abilities in reasoning and multimodality handling.
@@
The evolution of robotics datasets has demonstrated considerable diversity across various dimensions, particularly with the advancement of imitation learning and behavior cloning (BC). While many robotics benchmarks focus on evaluating model adaptability to new tasks, functionalities, or environments, there remains a gap in systematically evaluating different BC models at scale in both simulated and real-world settings. THE COLOSSEUM [35] addresses this gap by providing a systematic evaluation framework focused on robotic manipulation, evaluating generalization across 14 different environmental perturbations. Similar efforts include FactorWorld [45], which examines 11 variation factors across 19 tasks, and KitchenShift [46], which evaluates zero-shot generalization across 7 variation factors in kitchen environments.Several other specialized robotics benchmarks have emerged: RLBench [18] offers a suite of 100 manipulation tasks in simulation; RAVENS [16] focuses on vision-based manipulation; and FurnitureBench [14] provides reproducible real-world benchmarks for long-horizon complex manipulation. LIBERO [27] offers benchmarks for knowledge transfer in lifelong robot learning, while FMB [29] emphasizes generalizable robotic learning across complex tasks. Recent work has also introduced DUDE [40] for robotic document manipulation and ProcTHOR [10] for large-scale embodied AI using procedural generation.
@@ 3.1 Data
For version 0.1 of our benchmark, we utilize 53 of the 72 available OpenX datasets, as detailed in Table 3. We present results for 20 of these datasets for all 3 models, and have the full 53 for JAT.This subset was selected to ensure comprehensive coverage across different task types, embodiments, and environmental conditions while maintaining data quality and consistency. For datasets that did not include pre-defined evaluation sets, we have created and provided new evaluation splits to ensure robust assessment of model performance. The training splits of these 53 datasets comprise approximately 32 terabytes of data.
@@
The complete list of included datasets and their characteristics is provided in the appendix.
@@ 3.1.1 Dataset Curation
Our curation methodology consisted of several steps. First, we conducted a high-level review of dataset quality and accessibility, which resulted in the exclusion of three datasets: Austin BUDS, Austin Sailor, and Stanford Kuka Multimodal. For datasets that contained only training splits, we performed a detailed comparative analysis based on the robot platform used for data collection. This analysis considered multiple features: Robot model and morphology, Gripper specifications, Action space characteristics, Sensor configuration (number and type of RGB cameras, depth cameras, and wrist-mounted cameras), Presence of language annotations, Availability of camera calibration data, and Inclusion of proprioceptive information.
@@
Several additional datasets were excluded from version 0.1 of our benchmark due to technical limitations in their accessibility through the TensorFlow Datasets (TFDS) builder, which is the recommended data loading mechanism for OpenX. These compatibility issues will be addressed in future versions of the benchmark as the underlying infrastructure evolves. This careful curation process results in a benchmark that balances comprehensive coverage with practical considerations, ensuring that the included datasets provide meaningful evaluation scenarios while maintaining manageable computational requirements.
@@ 3.2 Models
In our evaluation, we focus on three recent vision-language-action (VLA) models that represent the current state-of-the-art in generalist robot learning: JAT (Jack of All Trades), GPT-4o, and OpenVLA. These models are particularly noteworthy for their ability to handle multiple modalities and their demonstrated capabilities across a wide variety of tasks.
@@
JAT emphasizes broad "generalist" multi-modal capabilities. GPT-4o is a powerful VLM, and allows for various approaches to map language output to action & control tasks. OpenVLA prioritizes open-source accessibility while maintaining competitive performance with larger closed-source models
@@ 3.3 Evaluation Metrics
We report several variations of MSE to provide a comprehensive performance analysis:
@@
    Average MSE (AMSE): Computed as the mean MSE across all timesteps in a dataset, AMSE enables direct comparison of model performance across different datasets and architectures.
@@
    Normalized AMSE (NAMSE): Calculated as an average of the min-max normalized MSEs over all the timesteps in the dataset, this metric normalizes prediction errors to each model’s error range, facilitating more equitable cross-dataset for a single model comparison by accounting for different scales in model outputs and dataset action spaces.
@@
    Completion Rate: We assess successful completion by comparing final predicted actions with ground truth final actions for all episodes in the dataset. While this serves as an approximate measure of task completion, it provides valuable insights into models’ ability to reach target states across trajectories.
The combination of these metrics allows us to evaluate both the fine-grained accuracy of action predictions and the overall task-completion capabilities of different models. This is particularly important in offline robotics, where environments and rewards are not available for policy evaluation.
@@ 4.1 Profiling Configuration
  • Multi-Image Processing: For timesteps with multiple available images, we select the primary image (typically designated with the keyword ‘image’).
@@
    • Continuous normalization: For $[0,1]$ to $[-1,1]$ conversion, we apply the formula: $y=2\cdot(x-orig_{low})/(orig_{high}-orig_{low})-1$. This was used by the authors in Kim et al. [22].
@@
Additional Considerations
We encountered cases where image observations were unavailable due to non-standard image key naming (e.g., ‘agentview_rgb’, ‘frontright_fisheye_image’) in some datasets. These were utilized for OpenVLA, but not the other models, as OpenVLA requires an image as part of its input. This specific case occurred with 2 datasets in particular, conq_hose_manipulation, and viola.
@@ 5 Results & Discussion
In our evaluation of vision-language-action models, we seek to answer several fundamental questions about their capabilities and limitations: (1) How do current VLM & VLA models perform across diverse robotics tasks and platforms, particularly in zero-shot settings? (2) What impact do different model architectures and training approaches (e.g., prompt engineering, robotics-specific training) have on performance? (3) How well do these models handle different action spaces and robot morphologies? (4) What are the current limitations and failure modes of these models in real-world robotics tasks? Through systematic analysis of three state-of-the-art models across 20 diverse datasets, we provide insights into these questions below.
Figure 2: AMSE values of GPT-4o, JAT, and OpenVLA Across 20 OpenX Datasets. JAT displays the poorest performance out of the 3 models with higher AMSE scores, while OpenVLA and GPT-4o demonstrate similar performance. OpenVLA displays consistent performance across most datasets
Table 1: Dataset Coverage and Action Space Characteristics
Dataset Name Registered Dataset Name In Pretraining OpenVLA Action Space Type
@@ 5.1 Average Model Performance Analysis
  • Complex manipulation tasks, particularly those involving large movements or multi-step sequences like Kitchen manipulation tasks.
  • Tasks requiring significant temporal reasoning or complex action sequences. This follows naturally as the models were assessed in a zero shot fashion.
@@ 5.1.2 Implications for Future Development
Figure 3: Normalized AMSE For GPT4o. GPT-4o demonstrates consistent NAMSE across all datasets, suggesting that the prompt engineering framework which provides detailed information about the action space, task, and environment, may be a key factor.
Figure 4: Normalized AMSE For JAT. JAT exhibits moderate variation in NAMSE across tasks, with a spike in the Stanford MaskVIT Dataset, while maintaining realtively consistent preformance for similar task types.
Figure 5: Normalized AMSE For OpenVLA. OpenVLA shows high variation in NAMSE values. As expected it displays very strong performance on tasks such as toto which is within its training distribution. OpenVLA shows a clear pattern of task-specific performance differences.
@@ 5.2.1 Model-Specific Performance Patterns
  • Exceptionally strong performance on pretrained tasks (e.g., toto with NAMSE: 0.003)
@@
  • Notable performance spike on stanford_mask_vit (NAMSE $\sim$0.57)
@@ 5.2.2 Cross-Model Insights
  • Complex manipulation tasks and multi-step operations consistently show higher NAMSE
@@
5.3 Key Takeaways
Our evaluation of VLM & VLA models reveals several fundamental insights about the current state of VLM & VLA models transferring to robotics tasks:
1. Prompt Engineering Impact: GPT-4o’s consistent performance across diverse tasks demonstrates that structured prompting with explicit action space information and environmental context may be as important as architectural choices. This suggests that future VLA development should consider incorporating structured task representations as a core design principle.
2. Specialization vs. Generalization: We observe a clear trade-off between specialized and general performance. OpenVLA shows superior performance on tasks within its training distribution but higher variation across tasks, while GPT-4o maintains more consistent but sometimes suboptimal performance. This highlights the ongoing challenge of developing models that can both specialize and generalize effectively.
3. Task Complexity Barriers: All models, regardless of architecture, struggle with complex manipulation tasks requiring multi-step planning or precise control. This consistent limitation suggests that current approaches may be missing key capabilities needed for complex robotics tasks.
4. Action Space Sensitivity: Performance varies significantly with different action space characteristics, particularly in tasks requiring precise control or complex action sequences. This suggests the need for more robust methods of handling diverse action spaces and robot morphologies.